# VMarker-Pro: Probabilistic 3D Human Mesh Estimation from Virtual Markers

### v3

@@ 4.6 Qualitative Results
VMarker. Figure 11 (top) presents some meshes estimated by our approach VMarker on natural images from the 3DPW test set. The rightmost case shows a typical failure where our method has a wrong pose estimate of the left arm due to heavy occlusion. We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates. As discussed in Sec. 15, we suppose that this is due to the advantages brought by the locality of the virtual marker representation, which can confine erroneous mesh estimates to a local range and thereby enhance robustness. Figure 2 (right) shows more examples where occlusion or truncation occurs, and our method can still get accurate or reasonable estimates robustly. Note that when truncation occurs, our method still guesses the positions of the truncated virtual markers.
Figure 11 (bottom) shows our estimated meshes on challenging cases, which indicates the strong generalization ability of our model on diverse postures and actions in natural scenes. Please refer to the supplementary for more quality results. Note that since the datasets do not provide supervision of head orientation, face expression, hands, or feet, the estimates of these parts are just in canonical poses inevitably.
VMarker-Pro. Figure 13 compares the visualization of VMarker and VMarker-Pro on the 3DPW [14] test set. We draw three hypotheses generated by VMarker-Pro, depicted in blue. Additionally, we include the results from the deterministic model VMarker (green). It is noticeable that, due to self-occlusion or object occlusion, VMarker might yield results that seem accurate in 2D projections yet are erroneous in 3D space, e.g.  the incorrect estimation of the right leg in the second case. This issue stems from the inherent ambiguity associated with occlusions. By leveraging the advantages of the data distribution learned by the denoising model, VMarker-Pro can generate more reasonable solutions, such as the naturally standing legs.
In Figure 14, we show more results of VMarker-Pro on unseen Internet images. It can be observed that VMarker-Pro possesses strong generalization capabilities, obtaining reasonable 3D estimations when faced with uncertain ambiguities, such as the various plausible poses of the runner’s right arm in the second case (as can be seen from the side view). Furthermore, even under complex actions and backgrounds, such as the person climbing in the third image, VMarker-Pro also demonstrates robust estimation results.
Figure 15 shows a typical failure case of VMarker-Pro on Internet image, where our method fails when heavy occlusion and ambiguity happen. Our method incorrectly estimates the right arm and right leg of the male dancer, which may be addressed using more powerful estimators or more diverse training datasets in the future.
@@ 5 Conclusion
In this paper, we present a novel intermediate representation virtual marker, which is more expressive than the prevailing skeleton representation and more accessible than physical markers. Our VMarker method, built on virtual markers, excels in detecting these markers in wild images to interpolate and reconstruct complete, realistically shaped meshes. To further address the ambiguity in 3D virtual marker prediction due to occlusions, we progress to VMarker-Pro, a probabilistic approach that generates multiple plausible meshes in alignment with the image. VMarker-Pro effectively captures the 3D virtual marker distribution through a conditional diffusion process, enabling precise modeling of 3D mesh distributions. Experiments demonstrate the effectiveness of our approaches and they show advanced generalization potential in spite of its simplicity.
Acknowledgments
This work was supported by National Science and Technology Major Project (2022ZD0114904).
Appendix A Post-processing on Virtual Markers
@@ B.1 Datasets
H3.6M [13]. Following previous works [6, 33, 16, 102], we use the SMPL parameters generated from MoSh [5], which are fitted to the 3D physical marker locations, to get the GT 3D mesh supervision. Following standard practice [6], we evaluate the quality of 3D pose of $14$ joints derived from the estimated mesh, i.e.  $\hat{\mathbf{M}}\mathcal{J}$. We report Mean Per Joint Position Error (MPJPE) and PA-MPJPE in millimeters (mm). The latter uses Procrustes algorithm [103] to align the estimates to GT poses before computing MPJPE. To evaluate mesh estimation results, we also report Mean Per Vertex Error (MPVE) which can be interpreted as MPJPE computed over the whole mesh.
@@
3DPW [14]. The 3D GT SMPL parameters are obtained by using the data from IMUs when collected. Following the previous works [87, 89, 47, 40], we use the train set of 3DPW to learn the model and evaluate on the test set. The same evaluation metrics as H3.6M are used. We further evaluate the performance of our probabilistic framework VMarker-Pro on two subsets of 3DPW, i.e.  3DPW-OC and 3DPW-PC [14, 95] which have both object and person occlusion. In this setting, we do not use the 3DPW training set for a fair comparison following [96].
@@
MPI-INF-3DHP [97] is a 3D pose dataset with 3D GT pose annotations. Since this dataset does not provide 3D mesh annotations, following [6, 33], we only enforce supervision on the 3D skeletons (Eq. 9) in mesh losses.
@@
UP-3D [15] is a wild 2D pose dataset with natural images. The 3D poses and meshes are obtained by SMPLify [29]. Due to the lack of GT 3D poses, the fitted meshes are not accurate. Therefore we only use the 2D annotations to train the 3D virtual marker estimation network as in [18].
@@
COCO [98] is a large wild 2D pose dataset with natural images. Previous work [17] used SMPLify-X [30] to obtain pseudo SMPL mesh annotations but they are not accurate. However, we find that if we project the 3D mesh to 2D image, the resulting 2D mesh vertices align well with the image. So we leverage the 2D annotations to train the virtual marker estimation network as in [18].
@@
SURREAL [27] is a large-scale synthetic dataset containing 6 million frames of synthetic humans. The images are photo-realistic renderings of people under large variations in shape, texture, viewpoint, and body pose. To ensure realism, the synthetic bodies are created using the SMPL body model, whose parameters are fit by the MoSh [5] given raw 3D physical marker data. All the images have a resolution of $320\times 240$. We use the same training split to train the model and evaluate the test split following [1].
B.2 Implementation Details
@@
Following common practice [6, 1, 17, 40, 16, 84, 89, 87], we conduct mix-training by using the above 2D and 3D datasets for experiments on the H3.6M and 3DPW datasets. To leverage the 3D pose estimation dataset, i.e.  MPI-INF-3DHP [97], we extend the $64$ virtual markers with the $17$ landmark joints (i.e.  skeleton) from the MPI-INF-3DHP dataset. For experiments on the SURREAL dataset, we use its training set alone as in [1, 36]. We implement the proposed method with PyTorch. The experiments for VMarker and VMarker-Pro are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs and 2 NVIDIA A100-PCIE-80G GPUs, respectively.
To train VMarker, the whole network is trained for 40 epochs with batch size 32 using Adam [100] optimizer. We set $\lambda_{3d},\lambda_{c},\lambda_{m},\lambda_{e}=100,0.1,0.1,0.1$ for each loss term. To train VMarker-Pro, the whole network is trained for 50 epochs with batch size 80 using Adam [100] optimizer. We set $\lambda_{diffusion}=1$ and keep the rest loss weights the same as training VMarker. For implementing LCN unit, we modify the original LCN network [77] by injecting a Cross-attention [82] module in each LCN layer. For more implementation details, please check https://github.com/ShirleyMaxx/VMarker-Pro.
@@ B.3 Additional Quantitative Results
Different loss terms. Table XII reports the MPVE error on H3.6M [13] test set when ablating different loss terms. The confidence loss [72] is used to encourage the interpretability of the heatmaps to have a maxima response at the GT position. Without the confidence loss, the error increases slightly. If ablating the surface losses, MPVE increases a lot, which demonstrates the smoothing effect of these two terms.
Comparison to fitting. In order to disentangle the ability of mesh regression from markers using $\hat{\mathbf{A}}$ with the ability to detect the virtual markers accurately from images, we first compute the estimation errors of the virtual markers. The MPJPE over all the virtual markers is $35.5$mm, which demonstrates that these virtual markers can be accurately detected from the images. We then fit the mesh model parameters to these virtual markers. Table XIII shows the metrics of the fitted mesh on the SURREAL [27] test set. As we can see, the fitted mesh has a similar error as our regression ones which uses the interpolation matrix $\hat{\mathbf{A}}$, which validates the accuracy of the estimated virtual markers.
@@ B.4 Additional Qualitative Results
Figure 19 shows more qualitative comparisons with Pose2Mesh [1] on the SURREAL test set in which has diverse body shapes. The skeleton representation used in Pose2Mesh [1] loses the body shape information so the method can only recover mean shapes, regardless of whether the person is slim or stout. This is caused by the limited skeleton representation bottleneck so that the model learns a mean shape for the whole training dataset implicitly. In contrast, both of our approaches VMarker and VMarker-Pro generate more accurate mesh estimation results with the help of virtual marker representation.
TABLE XII: MPVE errors on H3.6M [13] test set when ablating different loss terms.
VMarker w/o w/o w/o w/o
$\mathcal{L}_{conf}$ $\mathcal{L}_{pose}$ $\mathcal{L}_{normal}$ $\mathcal{L}_{edge}$
[…]
TABLE XIII: Results on SURREAL [27] test set when the mesh is obtained by fitting to the estimated virtual markers.
@@
VMarker 44.7 36.9 28.9
Figure 17: Meshes estimated by VMarker on Internet images with challenging cases (complex poses or extreme body shapes).
Figure 18: Meshes estimated by VMarker-Pro on Internet images with challenging cases (complex poses or extreme body shapes). For each case in a row, we present three hypotheses, showing both a projected view and a side view in each column. The rightmost column shows the overlapped side views to unveil the subtle differences.
@@
Figure 17 and 18 show more qualitative results of VMarker and VMarker-Pro on Internet images with challenging cases, respectively. These images include humans with diverse body shapes or complex poses. Our methods generalize well on the natural scenes. When occlusion happens, the probabilistic method VMarker-Pro could infer multiple reasonable poses, e.g.  the different angles of the left arm of the running woman in the first case in Figure 18.
Figure 19: Qualitative comparison between our methods and Pose2Mesh [1] on SURREAL test set [27]. Our approaches generate more accurate mesh estimation results on images of diverse body shapes. We show 2 hypotheses generated by VMarker-Pro.
Figure 20: Qualitative comparison between our method VMarker and Pose2Mesh [1], METRO [87] on 3DPW test set [14]. Our approach is more robust to occlusion and truncation and generates more accurate mesh estimation results that align images well.
Appendix C Human Subject Data

### v4

@@ 4.5 Qualitative Results
Figure 7 (top) presents some meshes estimated by our approach on natural images from the 3DPW test set. The rightmost case shows a typical failure where our method has a wrong pose estimate of the left leg due to heavy occlusion. We can see that the failure is constrained to the local region and the rest of the body still gets accurate estimates. We further analyze how inaccurate virtual markers would affect the mesh estimation, i.e. when part of human body is occluded or truncated. According to the finally learned coefficient matrix $\mathbf{\hat{A}}$ of our model, we highlight the relationship weights among virtual markers and all vertices in Figure 8. We can see that our model actually learns local and sparse dependency between each vertex and the virtual markers, e.g. for each vertex, the virtual markers that contribute the most are in a near range as shown in Figure 8 (b). Therefore, in inference, if a virtual marker has inaccurate position estimation due to occlusion or truncation, the dependent vertices may have inaccurate estimates, while the rest will be barely affected. Figure 2 (right) shows more examples where occlusion or truncation occurs, and our method can still get accurate or reasonable estimates robustly. Note that when truncation occurs, our method still guesses the positions of the truncated virtual markers.
Figure 8: (a) For each virtual marker (represented by a star), we highlight the top 30 most affected vertices (represented by a colored dot) based on average coefficient matrix $\mathbf{\hat{A}}$. (b) For each vertex (dot), we highlight the top 3 virtual markers (star) that contribute the most. We can see that the dependency has a strong locality which improves the robustness when some virtual markers cannot be accurately detected.
Figure 7 (bottom) shows our estimated meshes on challenging cases, which indicates the strong generalization ability of our model on diverse postures and actions in natural scenes. Please refer to the supplementary for more quality results. Note that since the datasets do not provide supervision of head orientation, face expression, hands, or feet, the estimates of these parts are just in canonical poses inevitably. Apart from that, most errors are due to inaccurate 3D virtual marker estimation which may be addressed using more powerful estimators or more diverse training datasets in the future.
@@ 5 Conclusion
In this paper, we present a novel intermediate representation Virtual Marker, which is more expressive than the prevailing skeleton representation and more accessible than physical markers. It can reconstruct 3D meshes more accurately and efficiently, especially in handling diverse body shapes. Besides, the coefficient matrix in the virtual marker representation encodes spatial relationships among mesh vertices which allows the method to implicitly explore structure priors of human body. It achieves better mesh estimation results than the state-of-the-art methods and shows advanced generalization potential in spite of its simplicity.
Acknowledgement
This work was supported by MOST-2022ZD0114900 and NSFC-62061136001.
Appendix
We elaborate on the post-processing implementation of the virtual markers and provide additional experimental details and results. At last, we discuss data from human subjects and the potential societal impact.
A.  Post-processing on Virtual Markers
@@ B.1  Datasets
Following previous works [19, 25, 26, 42], we use the SMPL parameters generated from MoSh [35], which are fitted to the 3D physical marker locations, to get the GT 3D mesh supervision. Following standard practice [19], we evaluate the quality of 3D pose of $14$ joints derived from the estimated mesh, i.e. $\hat{\mathbf{M}}\mathcal{J}$. We report Mean Per Joint Position Error (MPJPE) and PA-MPJPE in millimeters (mm). The latter uses Procrustes algorithm [13] to align the estimates to GT poses before computing MPJPE. To evaluate mesh estimation results, we also report Mean Per Vertex Error (MPVE) which can be interpreted as MPJPE computed over the whole mesh.
3DPW [54].
@@
The 3D GT SMPL parameters are obtained by using the data from IMUs when collected. Following the previous works [32, 33, 24, 61], we use the train set of 3DPW to learn the model and evaluate on the test set.
MPI-INF-3DHP [38]
@@
is a 3D pose dataset with 3D GT pose annotations. Since this dataset does not provide 3D mesh annotations, following [19, 25], we only enforce supervision on the 3D skeletons (Eq. (9)) in mesh losses.
UP-3D [28]
@@
is a wild 2D pose dataset with natural images. The 3D poses and meshes are obtained by SMPLify [2]. Due to the lack of GT 3D poses, the fitted meshes are not accurate. Therefore we only use the 2D annotations to train the 3D virtual marker estimation network as in [47].
COCO [34]
@@
is a large wild 2D pose dataset with natural images. Previous work [39] used SMPLify-X [41] to obtain pseudo SMPL mesh annotations but they are not accurate. However, we find that if we project the 3D mesh to 2D image, the resulting 2D mesh vertices align well with the image. So we leverage the 2D annotations to train the virtual marker estimation network as in [47].
SURREAL [53]
@@
is a large-scale synthetic dataset containing 6 million frames of synthetic humans. The images are photo-realistic renderings of people under large variations in shape, texture, viewpoint, and body pose. To ensure realism, the synthetic bodies are created using the SMPL body model, whose parameters are fit by the MoSh [35] given raw 3D physical marker data. All the images have a resolution of $320\times 240$. We use the same training split to train the model and evaluate the test split following [7].
B.2  Implementation Details and Computation Resource
@@
Following common practice [19, 7, 39, 61, 26, 23, 33, 32], we conduct mix-training by using the above 2D and 3D datasets for experiments on the H3.6M and 3DPW datasets. To leverage the 3D pose estimation dataset, i.e. MPI-INF-3DHP [38], we extend the $64$ virtual markers with the $17$ landmark joints (i.e. skeleton) from the MPI-INF-3DHP dataset. For experiments on the SURREAL dataset, we use its training set alone as in [7, 37]. We implement the proposed method with PyTorch. All the experiments are conducted on a Linux machine with 4 NVIDIA 16GB V100 GPUs. The whole network is trained for 40 epochs with batch size 32 using Adam [22] optimizer.
We evaluate the model complexity in terms of FLOPs (G) and the number of model parameters in Table 7. Compared to the most recent state-of-the-art methods that directly regress all mesh vertices, such as I2L-MeshNet [39], METRO [32], and Mesh Graphormer [33], our approach with virtual marker representation reduces the computation overhead by a large margin while getting better estimation quality. The last column shows the MPVE errors on 3DPW test set for performance reference.
[…]
I2L-MeshNet [39] ECCV’20 28.7 141.2 110.1
METRO [32] CVPR’21 153.0 397.5 88.2
Mesh Graphormer [33] ICCV’21 48.8 180.6 87.7
Ours 22.1 109.6 77.9
Table 7: Computation overhead comparison with the recent state-of-the-art methods that directly regress all 3D vertices. The rightmost column shows the MPVE errors on the 3DPW test set for performance reference.
Ours w/o $\mathcal{L}_{conf}$ w/o $\mathcal{L}_{pose}$ w/o $\mathcal{L}_{normal}$ w/o $\mathcal{L}_{edge}$
[…]
Table 8: MPVE errors on H3.6M [16] test set when ablating different loss terms.
[…]
None (Ours) 77.9 67.5 41.3
2 Arms 79.2 $\uparrow$ 1.3 68.2 $\uparrow$ 0.7 42.2 $\uparrow$ 0.9
2 Legs 78.3 $\uparrow$ 0.4 67.9 $\uparrow$ 0.4 41.7 $\uparrow$ 0.4
Body 78.6 $\uparrow$ 0.7 68.0 $\uparrow$ 0.5 41.8 $\uparrow$ 0.5
Random 78.7 $\uparrow$ 0.8 68.1 $\uparrow$ 0.6 41.9 $\uparrow$ 0.6
Table 9: Results on 3DPW [54] test set when different parts of virtual markers (VM) are occluded.
@@ B.3  Additional Quantitative Results
Figure 10: Meshes estimated by our approach on Internet images with challenging cases (complex poses or extreme body shapes).
Different loss terms.
Table 8 reports the MPVE error on H3.6M [16] test set when ablating different loss terms. The confidence loss [17] is used to encourage the interpretability of the heatmaps to have a maxima response at the GT position. Without the confidence loss, the error increases slightly. If ablating the surface losses, MPVE increases a lot, which demonstrates the smoothing effect of these two terms.
Robustness to occlusion.
We report results when different virtual markers are occluded by a synthetic mask in Table 9. The errors are slightly larger than the original image (None), which validates the effectiveness of the locality of the virtual marker representation. Occluding arm regions results in a larger error increase. This may be because the arm has larger variations in the dataset.
Comparison to fitting.
In order to disentangle the ability of mesh regression from markers using $\hat{\mathbf{A}}$ with the ability to detect the virtual markers accurately from images, we first compute the estimation errors of the virtual markers. The MPJPE over all the virtual markers is $35.5$mm, which demonstrates that these virtual markers can be accurately detected from the images. We then fit the mesh model parameters to these virtual markers. Table 10 shows the metrics of the fitted mesh on the SURREAL [53] test set. As we can see, the fitted mesh has a similar error as our regression ones which uses the interpolation matrix $\hat{\mathbf{A}}$, which validates the accuracy of the estimated virtual markers.
@@ B.4  Additional Qualitative Results
Figure 12 shows more qualitative comparisons with Pose2Mesh [7] on the SURREAL test set in which has diverse body shapes. The skeleton representation used in Pose2Mesh loses the body shape information so the method [7] can only recover mean shapes. For example, in Figure 12 (d) (e), the estimated meshes of Pose2Mesh tend to have the average body shape and fail to estimate the real body shape, regardless of whether the person is slim or stout. This is caused by the limited skeleton representation bottleneck so that the model learns a mean shape for the whole training dataset implicitly. In contrast, our approach with virtual marker representation generates more accurate mesh estimation results.
@@
     Ours      44.7      36.9      28.9
Table 10: Results on SURREAL [53] test set when the mesh is obtained by fitting to the estimated virtual markers.
@@
Figure 11: Typical failure cases. (a) The right arm has inaccurate shape estimation due to the inaccurate virtual marker estimation around the arm when occluded. (b) Our method treats the lower arm of another person as its own due to occlusion. (c) Interpenetration around the right hand.
Figure 14 shows more quality results of our approach on the 3DPW [54], H3.6M [16], MPI-INF-3DHP [38], and COCO [34] datasets. Figure 10 shows more qualitative results on Internet images with challenging cases, such as extreme body shapes or complex poses. Our method generalizes well on the natural scenes. Figure 11 shows typical failure cases, including inaccurate shape estimation and interpenetration, which are mainly caused by inaccurate 3D virtual marker estimation when occlusion occurs. But as expected, the rest body parts are barely affected due to the local and sparse property of the virtual marker.
Figure 12: Qualitative comparison between our method and Pose2Mesh [7] on SURREAL test set [53]. Our approach generates more accurate mesh estimation results on images of diverse body shapes.
Figure 13: Qualitative comparison between our method and Pose2Mesh [7], METRO [32] on 3DPW test set [54]. Our approach is more robust to occlusion and truncation and generates more accurate mesh estimation results that align images well.
Figure 14: Meshes estimated by our approach on images from the 3DPW [54] dataset (row 1-4), H3.6M [16] dataset (row 5), MPI-INF-3DHP [38] dataset (row 6), and COCO dataset (last 2 rows) [34].
C.  Human Subject Data